# Mechanistically Interpreting a Transformer-based 2-SAT Solver: An Axiomatic Approach

### v1

@@ 6 Conclusion
We presented a set of axioms that characterize a mechanistic interpretation and enable judgement of the interpretation’s validity with respect to the neural network under analysis. Using the axioms as a guide, we analyzed a Transformer-based model trained to solve the 2-SAT problem. Our axioms provide an automated and quantitative way of evaluating the quality of a mechanistic interpretation (via the $\epsilon$ values). Not only can the $\epsilon$’s serve as useful progress measures (in the sense of Nanda et al. [40]) to understand the training dynamics of models but can also help in the development of techniques for automated mechanistic interpretability analyses by serving as a useful evaluation metric. We intend to explore these directions next in conjunction with further development and formalization of the interpretability techniques used in this work.
Appendix A Axioms of Mechanistic Interpretation
@@
Axioms 5 and 6 are stated more informally. Unlike the earlier axioms requiring the observed behaviors of the mechanistic interpretation and the neural network to coincide, these axioms are concerned with the compilability of the mechanistic interpretation into a neural network. Both axioms assume the existence of a semantics-preserving compiler from $\lambda_{H}$ to $\lambda_{T}$ (denoted by $compiler^{\lambda_{H}}_{\lambda_{T}}$). While Axiom 5 requires the compiled version of $h$ to have the same structure and parameters as $t$ (i.e., syntactic equivalence), Axiom 6 requires the compiled version to only have the same structure as $t$. The requirements imposed by these last two axioms are hard to establish in practice, and we do not consider them for the analysis presented in this paper. We present these axioms as a goal for future work. Recent work on compilers from the RASP language [60] to Transformer models is a promising step in this direction [36].
@@
Appendix B More Details on Mechanistic Interpretability Analysis
B.1 First Layer as Parser
@@
The parsing behavior in the first layer cannot be fully described by token-to-token attention, and hence, we further decompose the standard QK-decomposition of Elhage et al. [22] to account for positional factors as well. In a Transformer, each token’s initial embedding is the sum of a positional embedding and a token embedding; hence, rather than viewing attention as from destination embeddings to source embeddings, we can equivalently express the pre-softmax scores as the sum of four sets of preferences—token-to-token attention, token-to-position attention, position-to-token attention, and position-to-position attention. In particular we can decompose the first-layer’s self-attention mechanism as follows. [^5] We use the QK circuit notation of Elhage et al. [22] Given token ${{t_{\text{src}}}}$ in position ${{p_{\text{src}}}}$ to token ${{t_{\text{dst}}}}$ in position ${{t_{\text{dst}}}}$, the pre-softmax score is
@@
Our observations in Figures 1 and 2 suggest that the tokens relevant to the final classification decision, are in position $24i+$ 2 for $100\leq i<$ 10, which, given the structure of the inputs to the model, are the second tokens of each clause of the form $x_{i}$ or $\neg x_{i}$ for $50\leq i<$ 5.
Figure 6: Decomponsed QK circuit of the first layer’s attention mechanism; the subset where the destination token is the second variable of a clause is shown.
If we restrict our view to destination tokens at this position, we can express the effects of the four components of the first layer’s QK circuit as shown in Figure 6. With the somewhat odd exception of the first clause, and, in particular, a first clause with second literal $x_{0}$ or $\neg x_{0}$, attention strongly de-prioritizes the parentheses but has weak preferences for the source token otherwise. Outside of destination tokens $x_{0}$ and $\neg x_{0}$, preferences for source position are fairly consistent, and hence, the core effect is a combination of position-to-position preferences (see the bottom right of Figure 6) and a preference against punctuation (the parentheses tokens ‘(’ and ‘)’). This can be seen in the token-to-position preferences in the bottom left, with an exception in the case of token 2, which may behave differently as all tokens to the left belong to the first clause, limiting the importance of learning more specific attention patterns. A general preference against punctuation can be seen in the token-to-position preferences (top left) as well. For the first several clauses, positional preferences encode a preference for the positions where parentheses occur, open parentheses in particular; however, the preference against parentheses shown in the token-to-position attention scores (corresponding to ${{W_{QK}^{\text{tok}\rightarrow\text{pos}}}}$) counteract that effect to result in a net effect of attention to the first literal of the clause.
@@
Now, we can perform similar analysis on the attention of the ‘:’ token. We observe in Figure 1 that the attention of the ‘:’ token in the first-layer is near uniform. To see why this occurs, see Figure 8, which fully describes the first-layer attention of ‘:’. These are derived from the four QK matrices using the known position $c=40$ of the token. The attention scores are consistently very small and have very little variation. As above, we can use this set of preferences to calculate a lower bound on the attention paid by the ‘:’ token to any individual token:
@@
Listing 1: First layer’s mechanistic interpretation in Python syntax.
Listing LABEL:fig:layer1_code shows our extracted mechanistic interpretation of the first layer as Python code.
B.2 Second Layer as Evaluator
Figure 9: Second-layer abstract attention preferences.

### v2

@@ 6 Conclusion
We presented a set of axioms that characterize a mechanistic interpretation and enable judgment of the interpretation’s validity with respect to the neural network under analysis. Using the axioms as a guide, we analyzed a Transformer-based model trained to solve the 2-SAT problem. We applied our axioms to validate the mechanistic interpretation of this 2-SAT model and a model trained to solve the modular addition task. Our axioms provide an automated and quantitative way of evaluating the quality of a mechanistic interpretation (via the $\epsilon$ values). Not only can the $\epsilon$’s serve as useful progress measures (in the sense of Nanda et al. (2023a)) to understand the training dynamics of models but can also help in the development of techniques for automated mechanistic interpretability analyses by serving as a useful evaluation metric.
Impact Statement
Acknowledgments
Nils Palumbo and Somesh Jha are partially supported by DARPA under agreement number 885000, NSF CCF-FMiTF-1836978 and ONR N00014-21-1-2492.
Appendix A More Related Work
Probing. Probing involves training a separate supervised classifier to predict human-understandable features of the data from the model’s internal activations (Alain & Bengio, 2017; Zou et al., 2023) and has found wide applications for understanding the representations learned by language models (Tenney et al., 2019; Li et al., 2023; Nanda et al., 2023b; Burns et al., 2023; Belrose et al., 2023; Huben et al., 2024; Gurnee et al., 2023; Wang et al., 2024). Even though a probe might be successful in predicting features from internal activations, it need not imply a causal relationship between these features and the model output (Belinkov, 2022). It has also been questioned whether the model activations genuinely represent the feature of interest or the probe itself learns these features by picking up on spurious correlations (Hewitt & Liang, 2019). However, recent works have demonstrated that for Transformer-based language models probes can be used to steer the model’s output behavior by manipulating the identified internal representations (Zou et al., 2023; Li et al., 2023).
Related to probing is the work on concept-based reasoning (Kim et al., 2018; Crabbé & van der Schaar, 2022; Yeh et al., 2021, 2020) that extracts representations of high-level concepts as geometric shapes in the neural network representation space (i.e., the activation spaces of inner layers of neural networks) and quantifies the impact of these high-level concepts on model outputs via attribution-based techniques.
Attention Pattern Analysis. There is a large body of work that attempts to understand the behavior of Transformer-based models by analyzing the attention patterns computed by the self-attention layer (Allen-Zhu & Li, 2023; Zhang et al., 2023; Ebrahimi et al., 2020; Vig & Belinkov, 2019; Hao et al., 2021; Qiang et al., 2022; Lu et al., 2021; Abnar & Zuidema, 2020; Chefer et al., 2021; Liu et al., 2023). However, there is an active ongoing debate about the validity of attention patterns as a tool for interpreting model behavior (Jain & Wallace, 2019; Pruthi et al., 2020; Bibal et al., 2022; Bastings & Filippova, 2020; Wiegreffe & Pinter, 2019).
Appendix B Axioms of Mechanistic Interpretation
@@
Axioms 5 and 6 are stated more informally. Unlike the earlier axioms requiring the observed behaviors of the mechanistic interpretation and the neural network to coincide, these axioms are concerned with the compilability of the mechanistic interpretation into a neural network. Both axioms assume the existence of a semantics-preserving compiler from $\lambda_{H}$ to $\lambda_{T}$ (denoted by $compiler^{\lambda_{H}}_{\lambda_{T}}$). While Axiom 5 requires the compiled version of $h$ to have the same structure and parameters as $t$ (i.e., syntactic equivalence), Axiom 6 requires the compiled version to only have the same structure as $t$. The requirements imposed by these last two axioms are hard to establish in practice, and we do not consider them for the analysis presented in this paper. We present these axioms as a goal for future work. Recent work on compilers from the RASP language (Weiss et al., 2021) to Transformer models is a promising step in this direction (Lindner et al., 2023).
@@
B.1 Extension to Circuits
While we focus on end-to-end analyses, our axioms are compatible with the analysis of subgraphs of a larger model.
B.1.1 Axioms for Arbitrary Computational Graphs
We can extend our axioms to operate directly on the graphs $G$ and $G^{\prime}$; we present a generalization of Axioms 1 through 4 below. Let $G=(V,E)$ be the concrete model expressed as a computational graph and let $G^{\prime}=(V^{\prime},E^{\prime})$ be the computational graph, isomorphic to $G$, representing the abstract model. Let the isomorphism between $G$ and $G^{\prime}$ be $\Pi$ and call the input and output vertices of $G$, $in$ and $out$, respectively; while this can be easily extended to multiple input and output vertices, we present the single input, single output case for simplicity. Let $execute(G)(x)$ return a mapping $val:V\rightarrow Val$ with the values of all intermediate nodes where $Val$ is the set of all possible values (i.e., vectors of reals) that the nodes can take. For a (partial) assignment to the vertices $assign:V^{\prime}\rightarrow Val$ with $\{in\}\subseteq V^{\prime}\subseteq V$, $propagate(G)(assign)$ returns a mapping of the same type $val:V\rightarrow Val$ which is the result of execution of $G$ where the provided intermediate assignments override the standard results of execution for those vertices. In this way, we can represent arbitrary interventions on $G$ and $G^{\prime}$.
We define an abstraction operator $\alpha_{v}$ for each concrete node $v\in V$ and a concretization operator $\gamma_{v}^{\prime}$ for each abstract node $v^{\prime}\in V^{\prime}$. $\alpha$ and $\gamma$ are the corresponding functions which apply the corresponding mappings to each vertex in the graph. Finally, define $select(v)(val)=val(v)$ for $val:V\rightarrow Val$ and $v\in V$; we overload and let $select(V^{\prime})(val)(v^{\prime})=val(v^{\prime})$ for $V^{\prime}\subseteq V$ and $v^{\prime}\in V^{\prime}$.
We can then define extensions of Axioms 1 to 4 below:
Axiom 7 ($\epsilon$-Prefix Equivalence).
$\forall v\in V.$
$\underset{x\sim\mathcal{D}}{Pr}[\alpha_{v}\circ select(v)\circ execute(G)(x)=select(\Pi(v))\circ execute(G^{\prime})\circ\alpha_{in}(x)]\geq 1-\epsilon$
Axiom 8 ($\epsilon$-Component Equivalence).
$\forall v\in V.$
$\underset{x\sim\mathcal{D}}{Pr}\left[\begin{array}[]{l}\alpha_{v}\circ select(v)\circ execute(G)(x)=\\ select(\Pi(v))\circ propagate(G^{\prime})\circ select(\Pi(p)\cup\{\Pi(in)\})\circ\alpha\circ execute(G)(x)\end{array}\right]\geq 1-\epsilon$
where $p=predecessors(v)$.
Axiom 9 ($\epsilon$-Prefix Replaceability).
$\forall v\in V.$
$\underset{x\sim\mathcal{D}}{Pr}\left[\begin{array}[]{l}select(out)\circ execute(G)(x)=\\ select(out)\circ propagate(G)\circ select(\{in,v\})\circ\gamma\circ execute(G^{\prime})\circ\alpha_{in}(x)\end{array}\right]\geq 1-\epsilon$
Axiom 10 ($\epsilon$-Component Replaceability).
$\forall v\in V.$
$\underset{x\sim\mathcal{D}}{Pr}\left[select(out)\circ execute(G)(x)=\left(\begin{array}[]{l}select(out)\circ propagate(G)\circ select(\{in,v\})\circ\\ \gamma\circ propagate(G^{\prime})\circ select(\Pi(p)\cup\{\Pi(in)\})\circ\\ \alpha\circ execute(G)(x)\end{array}\right)\right]\geq 1-\epsilon$
where $p=predecessors(v)$.
Note that our Axioms 1 to 4 are equivalent to these formulations, specialized to linear computational graphs.
In linear computational graphs, all dependencies on ancestor nodes are mediated by a node’s predecessor; hence, there is a limited need to consider parallel interventions, which involve simultaneously performing independent interventions on multiple nodes of the graph. In the case of prefix equivalence and prefix replaceability, any parallel interventions on linear graphs are equivalent to the intervention on the final node in the sequence and hence have no additional effect. While parallel extensions of component equivalence and component replaceability would strengthen evaluation in the linear setting, considering parallel interventions is particularly important in the nonlinear case.
Consider the following scenario: there are two sibling nodes in the graph that encode redundant computations and the concrete states output by both these nodes contain information needed by the downstream circuit. Also, let us say that the corresponding abstract model fails to faithfully capture the computation performed by either of these nodes. However, this failure may not be captured by the axioms without parallel interventions since the redundant copy of the computation in the sibling concrete node masks any intervention applied to the other concrete node.
One potential solution is to merge such redundant sibling nodes in the analyzed graph $G$, however, this prevents checking equivalence of the nodes independently. A more general solution is to extend Axioms 7 through 10 to accommodate arbitrary parallel interventions. We present a preliminary formulation of an axiom for parallel intervention below; a final formulation is left to future work. In particular, this axiom cannot be efficiently evaluated in its current form.
We first extend $execute$ and $propagate$ to accommodate execution of arbitrary mixtures of the graphs $G$ and $G^{\prime}$. These behave as before, except that any concrete inputs to an abstract operation are abstracted prior to execution, and, similarly, any abstract inputs to a concrete operation are concretized prior to evaluation. Let $interleave(G,G^{\prime},V^{\prime})$ be a graph identical to $G$ except that the operations for any nodes $v\in V^{\prime}$ are replaced with the operations for the corresponding nodes $\Pi(v)$ in $G^{\prime}$. Let $conditional\_abstract(V^{\prime})$ be a mapping on the values of the nodes $v\in V$ identical to $\alpha_{\Pi(v)}$ for $v\not\in V^{\prime}$ and the identity if $v\in V^{\prime}$. We can now express an axiom which generalizes Axioms 78, and 10 to parallel interventions:
Axiom 11 ($\epsilon$-Equivalence).
$\forall V^{\prime}\subseteq V,\forall v\in V$.
$\underset{x\sim\mathcal{D}}{Pr}\left[\begin{array}[]{l}select(v)\circ conditional\_abstract(V^{\prime})\circ execute(G)(x)=\\ select(v)\circ conditional\_abstract(V^{\prime})\circ execute(interleave(G,G^{\prime},V^{\prime}))(x)\end{array}\right]\geq 1-\epsilon$
The condition of Axiom 11 can be made even stronger by the treeification (Chan et al., 2022) of the graphs $G$ and $G^{\prime}$ prior to evaluation. This allows, for each node $v\in V$, independent selection of the set of ancestors whose operations are replaced with their abstract equivalents. With treeification, Axiom 11 can express prefix replaceability as well, leading it to generalize Axioms 7 through 10.
B.1.2 Linearizing Circuits
We can directly validate arbitrary circuits with the standard linear-graph axioms. In particular, we will describe how, for any computational graphs $G=(V,E)$ and $G^{\prime}=(V^{\prime},E^{\prime})$, which represent the concrete model $t$ and the abstract model $h$, respectively, we can construct a linearized computational graph. It is enough to show the construction for $G$.
We will do so by propagating partial assignments to the vertices $v\in V$ through a topologically sorted sequence of operations which compute the values of each node $v$ as a function of the values of the predecessors. First, we extend $\lambda_{T}$ to include multivariate functions $f(x_{1},x_{2},\dots,x_{n})$ for $x_{i}\in Var$. Suppose that $op:V\rightarrow\lambda_{T}$ returns the operation $op(v)$ computed by node each node $v$, and let $predecessors(v)$ return the predecessors of node $v$ in the argument order of $op(v)$. For simplicity, we assume that $G$ has exactly one input node $in$ and exactly one output node $out$.
Let $o:[|V|]\rightarrow V$ be a topological ordering of $G$. Let $idx:V\rightarrow[|V|]$, the inverse of $o$, return the position of each node $v$ in the topological ordering $o$. We again extend $\lambda_{T}$ to include the function $execute\_op(f,p,v)(env)$ which, given an $m$-ary function $f$, a node $v\in V$, a partial assignment to the vertices $env:V\rightarrow Val\cup\{\perp\}$ and a sequence of input nodes $p:[m]\rightarrow V$ such that $env(p(i))\in Val$ for each $1\leq i\leq m$, returns an updated environment $env^{\prime}$ of the same type, identical to $env$ except that $env^{\prime}(v)=f(env(p(1)),\ldots,env(p(m)))$. Finally, for $val\in Val$, let $input\_env(val)$ be a partial assignment to the vertices $input\_env(val):V\rightarrow Val\cup\{\perp\}$ which is $\perp$ everywhere but $in$, where we have $input\_env(val)(in)=val$.
Noting that $op(1)$ must always be $in$, we can define a decomposition of $t$ as follows:
$d_{t}=$ $\,\left(select(out)\circ execute\_op(op(o(|V|)),predecessors(o(|V|)),o(|V|))\right)$
[…]
$\,\circ execute\_op(op(o(2)),predecessors(o(2)),o(2))$
$\,\circ input\_env.$
Appendix C Remark on Extensional Equivalence
$x_{0}$(0)$x_{1}$(1)$x_{0}$(2)$x_{1}$(3)$1\over x_{0}$(4)$1\over x_{1}$(5)${1\over x_{0}}<{1\over x_{1}}$(6)(a)$x_{0}$(0)$x_{1}$(1)$x_{0}$(2)$x_{1}$(3)$x_{0}$(4)$x_{1}$(5)$x_{0}>x_{1}$(6)(b)
Figure 4: Extensionally equivalent mechanistic interpretations considered by Scheurer et al. (2023). Model (a) is the concrete model and the ground truth interpretation. Model (b) is the abstract model.
Scheurer et al. (2023) presents an example illustrating the extensional equivalence problem: in particular, they demonstrate that causal scrubbing cannot distinguish between two mechanistic interpretations with equivalent input-output behavior but which differ in internal behavior. See Figure 4 for the two models. We make copies of the $x_{0}$ and $x_{1}$ nodes in model (b) to ensure that the two models are isomorphic.
As observed by Scheurer et al. (2023), the two are indistinguishable to causal scrubbing, noting that we assume that $x_{0}$ and $x_{1}$ are both positive. This is a necessary consequence of the resampling technique used by causal scrubbing to derive its interventions. In particular, causal scrubbing replaces the values of concrete nodes with randomly selected values which are equivalent up to the abstract model.
Clearly, the concrete model is invariant with respect to resampling ablations derived from interpretation (a), which is identical to the concrete model. But, likewise, there is no effect from the corresponding intervention derived from interpretation (b). For instance, resampling concrete values of node (4), $1\over x_{0}$, up to equivalence in the abstract state of $x_{0}$ can have no effect. Hence, both interpretations are evaluated as perfect by causal scrubbing. For the same reason, these interpretations cannot be distinguished by interchange intervention accuracy (Geiger et al., 2022): interchange intervention accuracy replaces the values of intermediate nodes with those derived with different sets of inputs and evaluates the rate at which these interventions have an identical effects on the output of the abstract model on and the abstracted output of the concrete model.
If we restrict the set of allowable $\alpha$ and $\gamma$ functions, for example to linear mappings alone, Axioms 1 through 4 can distinguish between the two hypotheses. In particular, as no more than two points on the curve $(x,1/x)$ can be collinear, Axiom 1 cannot hold with $\epsilon=0$ on any distribution with more than four points in its support when we restrict $\alpha$ and $\gamma$ to be linear.
In the general case, with unrestricted $\alpha$ and $\gamma$, the interpretations are again indistinguishable. However to obtain this result, we must conduct meaningful computation in the abstraction and concretization functions, e.g. $\alpha_{4}(x)=\gamma_{4}(x)={1\over x}$; while not directly evaluated by the axioms, any analyst with a bias towards keeping complexity in the interpretation would prefer the ground-truth interpretation.
Appendix D More Details on Mechanistic Interpretability Analysis of the 2-SAT Model
Figure 5: Breakdown of the concrete model into concrete components, corresponding to the components $d_{t}[i]$.
The decomposition of the original Transformer 2-SAT solver is shown in Figure 5.
The abstraction operators from our analysis are shown in Listings LABEL:lst:alphas_1 and LABEL:lst:alphas_2 and the concretization operators are shown in Listings LABEL:lst:gammas_1 and LABEL:lst:gammas_2.
1alpha_0 = identity
2
3def clause_representation(literal_l, literal_r):
4 # Generate the input tensor for the formula with 10 copies of the clause
5 inputs = to_toks([Or(literal_l, literal_r)] * 10)
6
7 embeds = model.embed(inputs)
8 attn_out = embed + model.blocks[0].attention(
9 embeds,
10 mask=gen_attention_mask([4*i+2: clause_positions_and_parens(i)])
11 )
12 block_1_out = attn_out + model.blocks[0].mlp(attn_out)
13
14 # We derive the canonical representation by averaging across second literal positions
15 return mean(block_1_out[4*clause_idx+2] for clause_idx in range(10))
16
17clause_representations = {
18 Or(l,r): clause_representation(l, r)
19 for l in literals
20 for r in literals
21}
22
23def alpha_1(block_1_output):
24 second_literal_outputs = [block_1_outputs[4*i+2] for i in range(10)]
25 # Calculate cosine similarities between the actual representations output at
26 # the second variable positions and the canonical clause
27 # representations, select the clauses which best match
28 return [argmax_cosine_sims(out, clause_representations) for out in second_literal_outputs]
Listing 1: Abstraction operators for the first block in Python pseudocode.
1def alpha_2(residual_and_mlp_output, threshold=0.5):
2 mlp_output = residual_and_mlp_output[1]
3 return [mlp_output[i] >= threshold for i in evaluating_neurons]
4
5alpha_3 = identity
Listing 2: Abstraction operators for the second block in Python pseudocode.
1gamma_0 = identity
2
3# Mean output from the first transformer block, by position, on the training set
4mean_block_1_out = mean(model.intermediate_outputs("block_1", x) for x in train)
5
6# Constant on everything but second token positions
7def gamma_1(clauses):
8 output = mean_block_1_out.copy()
9 # Substitute the corresponding canonical clause representation
10 for i, clause in enumerate(clauses):
11 output[4*i+2] = clause_representations[clause]
12 return output
Listing 3: Concretization operators for the first block in Python pseudocode.
1# Mean residual from attention for the readout token on the training set
2mean_attn_out = mean(model.intermediate_outputs("attention_block_2", x)[-1] for x in train)
3
4# Output a constant residual term, and map predicted activations to
5# a large constant value
6def gamma_2(activation_model_outputs, high_activation=2):
7 mlp_out_concretization = torch.zeros(mlp_width)
8 for i, model_out in zip(evaluating_neurons, activation_model_outputs):
9 if model_out:
10 mlp_out_concretization[i] = high_activation
11
12 return mean_attn_out, mlp_out_concretization
13
14gamma_3 = identity
Listing 4: Concretization operators for the second block in Python pseudocode.
D.1 First Block as Parser
Figure 6: Average attention scores, grouped by source token position, for all heads, calculated over the test set. In the first block, we further average across destination token positions restricted to positions $24i+$ 2, where $100\leq i<$ 10 is the clause index. For the second block, we only consider the readout token as the destination.
@@ D.1 First Block as Parser
The parsing behavior in the first block cannot be fully described by token-to-token attention, and hence, we further decompose the standard QK-decomposition of Elhage et al. (2021) to account for positional factors as well. In a Transformer, each token’s initial embedding is the sum of a positional embedding and a token embedding; hence, rather than viewing attention as from destination embeddings to source embeddings, we can equivalently express the pre-softmax scores as the sum of four sets of preferences—token-to-token attention, token-to-position attention, position-to-token attention, and position-to-position attention. In particular we can decompose the first block’s self-attention mechanism as follows. [^5] We use the QK circuit notation of Elhage et al. (2021) Given token ${{t_{\text{src}}}}$ in position ${{p_{\text{src}}}}$ to token ${{t_{\text{dst}}}}$ in position ${{t_{\text{dst}}}}$, the pre-softmax score is
@@
Our observations in Figures 1 and 6 suggest that the tokens relevant to the final classification decision, are in position $24i+$ 2 for $100\leq i<$ 10, which, given the structure of the inputs to the model, are the second tokens of each clause of the form $x_{i}$ or $\neg x_{i}$ for $50\leq i<$ 5.
Figure 7: Decomposed QK circuit of the first block’s attention mechanism; the subset where the destination token is the second variable of a clause is shown.
If we restrict our view to destination tokens at this position, we can express the effects of the four components of the first block’s QK circuit as shown in Figure 7. With the somewhat odd exception of the first clause, and, in particular, a first clause with second literal $x_{0}$ or $\neg x_{0}$, attention strongly deprioritizes the parentheses but has weak preferences for the source token otherwise. Outside of destination tokens $x_{0}$ and $\neg x_{0}$, preferences for source position are fairly consistent, and hence, the core effect is a combination of position-to-position preferences (see the bottom right of Figure 7) and a preference against punctuation (the parentheses tokens ‘(’ and ‘)’). This can be seen in the token-to-position preferences in the bottom left, with an exception in the case of token 2, which may behave differently as all tokens to the left belong to the first clause, limiting the importance of learning more specific attention patterns. A general preference against punctuation can be seen in the token-to-position preferences (top left) as well. For the first several clauses, positional preferences encode a preference for the positions where parentheses occur, open parentheses in particular; however, the preference against parentheses shown in the token-to-position attention scores (corresponding to ${{W_{QK}^{\text{tok}\rightarrow\text{pos}}}}$) counteract that effect to result in a net effect of attention to the first literal of the clause.
@@
Figure 8: Softmax of expected attention scores by position for the second token of each clause, showing that for each second token position, attention on literal positions in the corresponding clause are expected to dominate, consistent with our interpretation of the first block’s attention mechanism as a parser.
@@
Now, we can perform similar analysis on the attention of the ‘:’ token. We observe in Figure 1 that the attention of the ‘:’ token in the first block is near uniform. To see why this occurs, see Figure 10, which fully describes the first-block attention of ‘:’. These are derived from the four QK matrices using the known position $c=40$ of the token. The attention scores are consistently very small and have very little variation. As above, we can use this set of preferences to calculate a lower bound on the attention paid by the ‘:’ token to any individual token:
@@
Listing 5: First component’s mechanistic interpretation in Python syntax.
Listing LABEL:fig:layer1_code shows our extracted mechanistic interpretation of the first component as Python code.
D.2 Second Block as Evaluator
Figure 11: Second-block abstract attention preferences.